Three pounds of perfect gas with R=38 ft.lb/lb.R and k= 1.667 have 300 Btu of heat added during a reversible nonflow constant pressure change of state. The initial temperature is 100 °F. Determine the

(A) final temperature

(B) change in heat ∆H

(C) W

(D) ∆U

(E) ∆S


1
Expert's answer
2022-02-10T15:09:17-0500

(1)


"C_p=\frac{kar}{k-1}=\frac{1.667\times38}{1.667-1}=\frac{94.97}{778}=0.1221BTU/lb R"

"Q=mc_p(T_2-T_1)"

"300=3\times0.1221(T_2-560)"

"T_2=1379=919\u00b0F"

(B)

"Q=mc_p(T_2-T_1)=H"

"H=300BTU"

(C)


"W=p(v_2-v_1)=p(\frac{mRT}{p_2}-\frac{mRT_1}{p_1})"

"W=pmR(\frac{T_2}{p}-\frac{T_1}{p})=mR(T_2-T_1)"

"W=3\times38\times(1379-560)=120.008BTU"

"(D)"


"C_v=\frac{R}{k-1}=\frac{38}{1.677-1}\times\frac{1}{778}=0.0732BTU/lb.R"

"\u2206U=mc_v(T_2-T_1)"


"∆U=3\times0.0732\times(1379-560)=179.85BTU"

(e)


"∆S=mc_pln\frac{T_2}{T_2}=4\times0.1221ln\frac{1379}{560}"

"\u2206S={0.3301}=BTU\/R"